Based on an equimolar mixture of water and ethanol and your knowledge of colligative properties (boiling point elevation), calculate the theoretical boiling point of your ethanol/water mixture. You will be distilling an equimolar mixture of ethanol and water. Based on its density at room temperature, calculate the correct volume of ethanol you will need to mix with 10.0 mL of water to make an equimolar mixture
Answer #1

The molarity of the described mixture is calculated:

m = (1 mol ethanol / 1 mol H2O) * (1 mol H2O / 18 g) * (1000 g / 1 Kg) = 55.6 m

The boiling temperature is calculated:

Tb = 100 + m * Kb = 100 + 55.6 * 0.513 = 156.11 ° C

The volume of ethanol needed is calculated:

V ethanol = 10 mL H2O * (0.997 g / 1 mL) * (1 mole H2O / 18 g) * (1 mole ethanol / 1 mole H2O) * (46.07 g ethanol / 1 mole) * (1 mL ethanol / 0.789 g ) = 32.34 mL

If you liked the answer, please rate it in a positive way, you would help me a lot, thank you.